OpenStudy (anonymous):

Use L'hopital's rule to evaluate the limit: lim x->infinity (1+(pi)sqrt(2)/x)^x

OpenStudy (solomonzelman):

\(\large\color{#000000}{\displaystyle\lim_{x \rightarrow ~\infty }\left(1+\frac{\pi\sqrt{2}}{x}\right)^x}\) like this?

OpenStudy (solomonzelman):

If that is the case, that I will give you the following limit property: \(\Large\color{#11aa11}{\displaystyle\lim_{x \rightarrow ~\infty }\left(1+\frac{\rm a}{x}\right)^x=e^{\rm a} }\)

OpenStudy (solomonzelman):

and that is regardless of the value of \(\color{#11aa11}{\rm a }\).

OpenStudy (anonymous):

I did come to that as well but i am confused as to why they asked me to use l'hopital rule

OpenStudy (solomonzelman):

If I did interpret the limit correctly, then the best I can do is to use Striling's approximation to show this (and then to attempt to use the squeeze thrm to show the rule for nonintegers)

OpenStudy (solomonzelman):

If this is the limit (as I posted), then I doubt\(^\infty\) about L'Hospital's rule...

OpenStudy (anonymous):

I was thinking the same

OpenStudy (anonymous):

so would i just substitute the (pi)sqrt(2) as a?

OpenStudy (solomonzelman):

Well, yes...

OpenStudy (anonymous):

ok thank you

OpenStudy (solomonzelman):

Yw`

OpenStudy (irishboy123):

i reckon you can use l'H, but it's a mess \(\large \lim_{x \rightarrow ~\infty } \left( 1+\frac{ a}{x}\right)^x\) \(\large = \lim_{x \rightarrow ~\infty } exp\left\{ \ln \left( 1+\frac{ a}{x}\right)^x \right\} \) the hand wavey bit next... \(\large = exp \left\{ \lim_{x \rightarrow ~\infty } \ln \left( 1+\frac{ a}{x}\right)^x \right\}\) \(\large = exp \left\{ \lim_{x \rightarrow ~\infty } x \, \ln \left( 1+\frac{ a}{x}\right) \right\}\) \(\large = exp \left\{ \lim_{x \rightarrow ~\infty } \dfrac{ \ln \left( 1+\frac{ a}{x}\right) }{1/x} \right\}\) so that's a \(\dfrac{0}{0}\), game on! because the exp function is so cool, like in laplace transforms, i reckon you can run with this and get something like solomon's answer.

OpenStudy (solomonzelman):

yes, indeed, very clever! When I differentiated on top and bottom, and simplified, I get, \(\large\color{#000000}{ \displaystyle\lim_{x \rightarrow ~\infty}(e)^{^{\displaystyle{\tiny~}ax/(x+a) }} }\) \(\large\color{#000000}{ \displaystyle u=x+a }\) \(\large\color{#000000}{ \displaystyle u-a=x }\) \(\large\color{#000000}{ \displaystyle ax=a(u-a)}\) And if x approaches infinity, so does the u, so... \(\large\color{#000000}{ \displaystyle\lim_{u \rightarrow ~\infty}(e)^{^{\displaystyle{\tiny~}a(u-a)/u }} }\) \(\large\color{#000000}{ \displaystyle\lim_{u \rightarrow ~\infty}(e)^{^{\displaystyle{\tiny~}(au-a)/u }} }\) \(\large\color{#000000}{ \displaystyle\lim_{u \rightarrow ~\infty}(e)^{^{\displaystyle{\tiny~}a-(a/u) }} }\) \(\large\color{#000000}{ \displaystyle\lim_{u \rightarrow ~\infty}(e)^{^{\displaystyle{\tiny~}a-0}} }\) \(\large\color{#000000}{ \displaystyle\lim_{u \rightarrow ~\infty}(e)^{^{\displaystyle{\tiny~}a}} \color{blue}{\bf =e^a }}\)